OpenStudy (anonymous):

PLEASE PLEASE HELP The consumer price index (CPI) indicates the relative change in price over time for a fixed basket of goods and service. In general if the rat of inflation averages r% over n years, then the CPI after n years is CPI=CPI (base 0) (1+(r/100))^n, where CPI(base 0) is the CPI index at the beginning of the n-years. (a) the CPI was 151.5 for 1995 and 202.4 for 2003. Assuming that annual inflation remained constant for this time period, determine the average annual inflation rate. a) the average annual inflation rate from 1995 to 2003 was approximately__% (b) Using the inflation rate from part a in what year will the CPI reach 256? b) the CPI will reach 256 in the year ________

OpenStudy (anonymous):

\[C=C_0\left(1+{r\over100}\right)^n\]

OpenStudy (anonymous):

a) C0=151.5 C=202.4 , n = 2003-1995=8 find "r"

OpenStudy (anonymous):

so it will look like\[202.4=151.5\left( 1+\frac{ r }{ 100 } \right)^{2003-1995}\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

I got a negative number its -83.3 which i think is wrong?

OpenStudy (anonymous):

how do you get r by itself since i think I am doing wrong?

OpenStudy (anonymous):

first divide both sides by 151.5 then take logarithms to bring the exponents down

OpenStudy (anonymous):

then simplify then take inverse logarithm to get the "r" out

OpenStudy (anonymous):

can you refresh me on how to take the logarithms to bring the exponent

OpenStudy (anonymous):

enter the simplified form first.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so i did 202.4/151.5

OpenStudy (anonymous):

k..

OpenStudy (anonymous):

so how will you get rid of the exponent?

OpenStudy (anonymous):

by using log

OpenStudy (anonymous):

\[\Large{\log(a^m)=m\times\log(a)}\]

OpenStudy (anonymous):

so it will be log(202.4/151.5)=8log(1+(r/100))

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

simplify it to \[\log(\ldots)=\ldots\]

OpenStudy (anonymous):

okay so would it look like log202.4)-log(151.5)=8log(1+(r/100)) and thats all I could think of

OpenStudy (anonymous):

you can use a calc to get the numbers...

OpenStudy (anonymous):

but I will you put them all to one side and r by itself? from that

OpenStudy (anonymous):

yes. you'd get a simple form like I showed above..

OpenStudy (anonymous):

log of 1+r/100 on one side and a number on the other side..

OpenStudy (anonymous):

so it will look like log of 1 (r/100) =.125797

OpenStudy (anonymous):

@electrokid

OpenStudy (anonymous):

good. now, I presume that is "log base 10" then \[1+{r\over100}=10^{\rm number}\]

OpenStudy (anonymous):

10^number you got on the right side

OpenStudy (anonymous):

okay?

OpenStudy (anonymous):

can you show me how it will look like

OpenStudy (anonymous):

just did.. look two steps above. :)

OpenStudy (anonymous):

remove the "log" on the left, and raise "10" to the power of that number

OpenStudy (anonymous):

so it will look like 1+(r/100)=10^.25797

OpenStudy (anonymous):

exactly.. solve for "r" like a good old algebra problem.

OpenStudy (anonymous):

i got 33.60%

OpenStudy (anonymous):

\[{1\over8}\log\left(202.4\over151.5\right)=\log\left(1+{r\over100}\right)\]

OpenStudy (anonymous):

re do the left part

OpenStudy (anonymous):

APR values are usually < 10%

OpenStudy (anonymous):

i got 3.69

OpenStudy (anonymous):

how will you do part b?

OpenStudy (anonymous):

good.

OpenStudy (anonymous):

use this "r" the C0 = C for 1998 C = 256, find "n"

OpenStudy (anonymous):

similar steps.. no need to take 10^ just the log will do the "n" is the number of years after 1998. so, find what year it'd be

OpenStudy (anonymous):

i keep getting a neg. number

OpenStudy (anonymous):

@electrokid

OpenStudy (anonymous):

follow the same steps should get it

OpenStudy (anonymous):

\[256=151.5\left(1+{3.7\over100}\right)^n\]

OpenStudy (anonymous):

okay so i did log(256)-log151)=nlog(1+.037)

OpenStudy (anonymous):

how will you get n buy itself from that? @electrokid

OpenStudy (anonymous):

is there a chance you can show me how it will look like if n was by itself?

OpenStudy (anonymous):

\[\log\left(256\over151\right)=n\times\log(1+0.037)\]

OpenStudy (anonymous):

yes now how do you just get n by itself. so just divided?

OpenStudy (anonymous):

yes.. use a calculator to gt the logs and divide.

OpenStudy (anonymous):

i got 70

OpenStudy (anonymous):

but wouold the 256 be 202.4 instead and 252 be 151.5

OpenStudy (anonymous):

no. it'd be 151.5 the answer is n=14.5 (approx)

OpenStudy (anonymous):

then where did the 256 come from again?

OpenStudy (anonymous):

check with the quesrtion

OpenStudy (anonymous):

oh okay thank you! sorry was confused!!! thank you for helping me while i struggle!

OpenStudy (anonymous):

:) you are welcome.